OpenStudy (anonymous):

The cost of running a small factory is modeled by the function where x is the number of routers (in hundreds) sold weekly and C is in dollars. C(x)= x^3-3x^2+18x The company sells routers for 42 dollars each. Write the equation for revenue as a function of number sold. (Hint: revenue = number sold * price.) R(x) =

OpenStudy (anonymous):

I fail to see the complexity of this problem. 'revenue = number sold * price' revenue R(x) number sold 'x is the number of routers (in hundreds) sold weekly' price 'for 42 dollars each' therefore it should be simply 42*x?

OpenStudy (anonymous):

You are so right, I forgot to attach the bottom half of the problem, which i had trouble with. The rest of the question is 1. Find the number of routers sold which will maximize weekly profit. (Hint: profit = revenue minus cost). 2. What is the company's maximum profit? I appreciate the help!

OpenStudy (anonymous):

Let profit be a function of numbers sold: P(x) P(x) = R(x) - C(x) P(x) = 42x - x^3 + 3x^2 - 18x P(x) = -x^3 + 3x^2 + 24x Now find when P'(x) = 0; Is when P(x) at extrema 0 = -3x^2 + 6x + 24 ... x = 4 or -2 now find 2nd order derivatives at these 2 values to determine nature of extrema P''(x) = -6x + 6 P''(4) = -18 < 0 -2 is an unrealistic solution to the problem, therefore there is no need to examine it further Now at weekly production of 4 the weekly profit is maximized profit at that point: P(4) = -4^3 + 3*4^2 + 24*4 = -64 + 48 + 96 = 80

OpenStudy (anonymous):

thank you so much!